OpenStudy (anonymous):

10w + 4w = -84 please show all work

OpenStudy (anonymous):

10w + 4w = -84 14w=-84 14w/14=-84/14 w=-6

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem